Striking a match stick on the matchbox involves the reaction of potassium chlorate and red phosphorus with the following equation: 10KClO3+3P4=3P4O10+10KCl. How much of each reactant will be needed to produce 105.8g pf tetraphosphorus decaoxide (P4O10)?


1
Expert's answer
2021-02-23T04:32:16-0500

1) n(P4) = n(P4O10);

m(P4)/M(P4) = m(P4O10)/M(P4O10);

m(P4) = M(P4)*m(P4O10)/M(P4O10) = 123.90*105.8/283.89 = 46.175 (grams)

2) Analogously, 10n(P4O10) = 3n(KClO3)

m (KClO3) = 10M(KClO3)*m(P4O10)/3M(P4O10) = 10*122.55*105.8/283.89/3 = 149 (grams).